OpenStudy (anonymous):

volume of a cylinder is1131.0 cubic centimeters. the lengh of the cylinder is 40 centimeters. what is the radius of the cylinder

OpenStudy (anonymous):

pi r^2 (40) = 1131 r =sqrt (1131/(40pi))

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

helloooooooooooooooooooo

OpenStudy (anonymous):

hello are you still there

myininaya (myininaya):

volume of cylinder is pi* r^2*h

myininaya (myininaya):

V=1131.0, h=40 solve for r

OpenStudy (anonymous):

what does ^ mean

myininaya (myininaya):

it means to the power of

OpenStudy (anonymous):

how do i solve for r

OpenStudy (anonymous):

okay thanx

myininaya (myininaya):

use the equation to solve for r: you have V=pi*r^2*h were h=40 and V=1131

myininaya (myininaya):

1131=pi*r^2*40

OpenStudy (anonymous):

i like mtininaya better please help me

myininaya (myininaya):

so since we want to solve for r, we need to divide both sides by 40 and pi

myininaya (myininaya):

1131/(pi*40)=r^2

myininaya (myininaya):

take the square root of both sides to get r

OpenStudy (anonymous):

im confused sorry can u go step by step

myininaya (myininaya):

click on that up there

myininaya (myininaya):

hows that?

myininaya (myininaya):

I divided both sides by h and pi to get r^2 by itself. i took square root of boths to get r. and then volume=V=1131 and height=h=40 and radius=r

OpenStudy (anonymous):

is it 1 centimeter

OpenStudy (anonymous):

is it 1 centimeter

myininaya (myininaya):

did you look at the attachment?

myininaya (myininaya):

the radius is sqrt(1131/[40pi])

OpenStudy (anonymous):

so to divide 1131/40pi i first mulyiply 40 times 3.14 right?

myininaya (myininaya):

If you do that, you will get an approximation. I would stop where I stoped

OpenStudy (anonymous):

bt i dnt get what the answer is because in my paper is multipal choice. a.1 centimeter b.2 centimeter c.3centimeter d.centimeter

myininaya (myininaya):

ok so they rounded it?

myininaya (myininaya):

i get 3 centimeters then

OpenStudy (anonymous):

can u show me how u got everything in the attach file

myininaya (myininaya):

what kindof calculator do you have?

myininaya (myininaya):

ti83?

OpenStudy (anonymous):

the one in the computer

myininaya (myininaya):

ok so type 40

OpenStudy (anonymous):

40

myininaya (myininaya):

hit the multplication sign

myininaya (myininaya):

type 3.14

myininaya (myininaya):

hit the equal

myininaya (myininaya):

remember 125.6

myininaya (myininaya):

now do 1131 and hit the division sign then type in 125.6

myininaya (myininaya):

then last step hit the square root sign